# The linearization approach to the Calderón problem revisited: reconstruction via the Born approximation

### v1

@@ Abstract.
Linearization techniques are widely used in the analysis and numerical solution of the Calderón inverse problem, even if their theoretical basis is not fully understood. In this article, we study the effectiveness of linearization for reconstructing a conductivity from its Dirichlet-to-Neumann (DtN) map, combining rigorous analysis with numerical experiments. In particular, we prove that any DtN map arising from a radial conductivity in the unit ball of $\mathbb{R}^{d}$ admits an exact representation as a linearized DtN map for a uniquely determined integrable function, the Born approximation. We linearize on a family of background conductivities that includes the constant case, giving a rigorous foundation for linearization-based methods in this framework. We also characterize the Born approximation as a solution of a generalized moment problem. Since this moment problem is formally well-defined even for non-radial conductivities, we use it to develop a numerical algorithm to reconstruct the Born approximation of a general conductivity on the unit disk. We provide numerical experiments to test the resolution and robustness of the Born approximation in different situations. Finally, we show how it can be used as the starting point of an algorithm for reconstructing a conductivity from its DtN map.
@@ 1.1. Setting and results
where $\gamma$ varies in some class of admissible conductivities. This map is differentiable, and the inversion of its differential $\mathrm{d}\Phi_{\sigma}$ at some reference conductivity $\sigma$ is known as the linearized Calderón problem. Since the work of Calderón [9], understanding this linearized version of the inverse problem has been a cornerstone in the strategy for solving the full nonlinear problem from both theoretical and numerical viewpoints.
Calderón’s inverse problem has been thoroughly investigated, partly due to its connection to a number of imaging techniques, such as Electrical Impedance Tomography (EIT). A wide range of linearization approaches to numerical reconstruction in EIT involve determining a function $\gamma^{\mathrm{B}}_{\sigma}$ that solves
@@
where $\gamma$ is the conductivity one desires to approximate with respect to a background conductivity $\sigma$ (or a sufficiently close prior); see [25] for instance. This is merely formal, since it is not clear that such a function exists. Motivated by the linear approximation in inverse scattering theory developed by Born in the 1920s, following [4, 5], we denote $\gamma^{\mathrm{B}}_{\sigma}$ as the Born approximation of $\gamma$ with respect to $\sigma$.
Though widely used in EIT, there are very few rigorous results on the existence of $\gamma^{\mathrm{B}}_{\sigma}$. The question of existence is subtle, and it is closely connected with the problem of characterizing the set of DtN maps. To the best of our knowledge, the existence of $\gamma^{\mathrm{B}}_{\sigma}$ has been rigorously established only when $\sigma=1$, $\Omega=\mathbb{B}^{d}\subset\mathbb{R}^{d}$ is the unit ball, and $\gamma\in W^{2,p}(\mathbb{B}^{d})$ is radial; see [17, Theorem 1] [^1] That result proves the existence of the Born approximation as the solution of a certain moment problem involving the spectrum of the DtN map. The connection with the Fréchet differential of $\Phi$ is presented in the proof of 1. (see also [18, 32] for rigorous results on the existence of the Born approximation for DtN maps associated with Schrödinger operators).
If $\gamma^{\mathrm{B}}_{\sigma}$ exists, then the strong injectivity properties of $\mathrm{d}\Phi_{\sigma}$ imply that one has a two-step factorization of the inverse problem: a linear step that asks for reconstructing the Born approximation $\gamma^{\mathrm{B}}_{\sigma}$ from the DtN map, and a non-linear one that consists of the inversion of the map $\Phi_{\sigma}^{\mathrm{B}}$ that sends the function $\gamma$ to the function $\gamma^{\mathrm{B}}_{\sigma}$. More precisely, using that $\Lambda_{\sigma}=\mathrm{d}\Phi_{\sigma}(\sigma)$ (see Proposition 2.1), (1.4) becomes
@@
This type of factorization can be formulated in a general context and has been applied under different names to solve a number of ill-posed inverse problems (see the survey [31]). In [17] it is shown that when $\gamma\in W^{2,p}(\mathbb{B}^{d})$ is a radial conductivity, $\gamma^{\mathrm{B}}_{1}$ exists and uniquely determines $\gamma$. Moreover, in this case, the map $(\Phi_{1}^{\mathrm{B}})^{-1}:\gamma^{\mathrm{B}}_{1}\longmapsto\gamma$ is Hölder continuous under suitable a priori assumptions on the conductivities (that unlike most conditional stability results, do not imply any compactness on the admissible set of conductivities). In other words, the well-known fact that $\Phi^{-1}:\Lambda_{\gamma}\longmapsto\gamma$ is discontinuous (i.e., the inverse problem is ill-posed) is due to the linear part of the factorization: the inversion of $\mathrm{d}\Phi_{\sigma}$. We say that the inverse problem admits a stable factorization if $\Phi$ can be factorized into a linear ill-posed map and a non-linear stable map. In addition to the results in [17] for the conductivity case, this property has also been proven in [18] for the radial Schrödinger case and an associated inverse spectral problem, and locally in [32] for the radial fixed energy Schrödinger case.
The objective of this article is twofold. First, we present a systematic approach to establishing the existence of the Born approximation; as a byproduct of our analysis, we broaden the class of conductivities $\sigma$ and $\gamma$ for which (1.4) can be rigorously solved. More specifically, we show that $\gamma^{\mathrm{B}}_{\sigma}$ exists as an integrable function as soon as $\Omega=\mathbb{B}^{d}$ for $d\geq 2$, $\gamma$ is a $W^{2,\infty}$ radial conductivity, and $\sigma$ is
@@
We denote by $\mathcal{E}^{\prime}(\Omega)$ the space of distributions with compact support contained in $\Omega$. It is shown in Remark 2.2 that, for $\sigma\in\mathcal{C}^{\infty}(\overline{\Omega})$, one can extend $\mathrm{d}\Phi_{\sigma}$ from its natural domain $L^{\infty}(\Omega)$ to $L^{\infty}(\Omega)+\mathcal{E}^{\prime}(\Omega)$.
@@
Suppose that $\Omega:=\mathbb{B}^{d}$ is the unit ball in $\mathbb{R}^{d}$, that $\gamma\in W^{2,\infty}(\mathbb{B}^{d})$ is a radial conductivity, and that $\kappa\in(-\infty,\lambda_{\nu_{d},1}^{2})$. Then there exists a unique distribution $\gamma^{\mathrm{B}}_{\sigma_{\kappa,d}}\in W^{2,\infty}(\mathbb{B}^{d})+\mathcal{E}^{\prime}(\mathbb{B}^{d})$ such that:
(1.8) $\mathrm{d}\Phi_{\sigma_{\kappa,d}}(\gamma^{\mathrm{B}}_{\sigma_{\kappa,d}})=\Lambda_{\gamma}.$
In addition, $\gamma^{\mathrm{B}}_{\sigma_{\kappa,d}}$ is an integrable function that satisfies:
@@
1 generalizes to $\kappa\neq 0$ some of the results in [17]. The uniqueness statement actually holds in a slightly more general setting, see Theorems 2.7 and 2.8. We also investigate the structure of the Born approximation for non-radial conductivities in the unit disk when $d=2$. In particular, we observe that the Born approximation can be characterized as the solution to a certain complex moment problem, and we obtain explicit formulas to compute these moments.
This leads us to the second objective. As a byproduct of our analysis of the Born approximation via the moment problem, we provide a simple and general framework for the numerical inversion of $\mathrm{d}\Phi_{\sigma}$. This framework is robust enough to hold for any choice of $\sigma$, or if Neumann-to-Dirichlet data is used instead of the DtN map. We apply it to a series of numerical experiments that are described below.
The precise moment problem we solve is described as follows. Let $(e_{m})_{m\in\mathbb{N}}$ be an orthonormal basis of $L^{2}(\partial\Omega)$ such that $e_{m}\in H^{1/2}(\partial\Omega)$ for all $m\in\mathbb{N}$. Define for $F\in L^{\infty}(\Omega)$,
$\mathfrak{m}_{\ell,m}^{\sigma}[F]:=\left\langle\overline{e_{\ell}},\mathrm{d}\Phi_{\sigma}(F)e_{m}\right\rangle_{H^{1/2}\times H^{-1/2}},\qquad\ell,m\in\mathbb{N}.$
By direct computation of the Fréchet derivative $\mathrm{d}\Phi_{\sigma}$ (see Proposition 2.1), one can prove that
$\mathfrak{m}_{\ell,m}^{\sigma}[F]=\int_{\Omega}F(x)\overline{\nabla u_{\ell}^{\sigma}}\cdot\nabla u_{m}^{\sigma}\,dx,$
where $u_{m}^{\sigma}$ is the solution of (1.1) with $f=e_{m}$. Since (1.5) implies that
$\mathfrak{m}_{\ell,m}^{\sigma}[\gamma^{\mathrm{B}}_{\sigma}]=\left\langle\overline{e_{\ell}},\Lambda_{\gamma}e_{m}\right\rangle_{H^{1/2}\times H^{-1/2}},\qquad\ell,m\in\mathbb{N},$
in order to compute $\gamma^{\mathrm{B}}_{\sigma}$, one needs to solve the generalized moment problem
@@
(1.9) $\int_{\Omega}\gamma^{\mathrm{B}}_{\sigma}\,\overline{\nabla u_{\ell}^{\sigma}}\cdot\nabla u_{m}^{\sigma}\,dx=\left\langle\overline{e_{\ell}},\Lambda_{\gamma}e_{m}\right\rangle_{H^{1/2}\times H^{-1/2}},\qquad\ell,m\in\mathbb{N},$
where the matrix elements of the DtN map in the RHS are assumed to be known from boundary measurements. Clearly, this method requires a good knowledge of the solutions $(u_{m}^{\sigma})_{m\in\mathbb{N}}$ of the conductivity equation for $\sigma$.
The previous moment problem is an ill-posed problem that can be numerically solved as a least-squares problem with regularization. It is worth noting that, in some situations, solutions of (1.9) are uniquely determined if one takes moments with respect to an orthonormal basis $(e_{m})_{m\in\mathbb{N}}$ of a proper closed subspace of $L^{2}(\partial\Omega)$, which suggests that (1.9) is overdetermined (this is indeed the case when $\kappa=0$ and $\Omega=\mathbb{B}^{2}$, see Remark 3.7).
@@
Finally, we give numerical experiments illustrating the use of (1.9) in several examples of non-radial $\gamma$ for $\sigma=\sigma_{\kappa,2}$ as defined by (1.7) on the unit disk $\Omega=\mathbb{B}^{2}=\mathbb{D}$ in $\mathbb{R}^{2}$. Moments are taken with respect to the basis $(e_{m})_{m\in\mathbb{N}}$ of spherical harmonics in $\mathbb{S}^{1}$. This choice of $\sigma$ has the advantage of providing good a priori knowledge of the corresponding solutions $u_{m}^{\sigma}$ and encompasses the framework in which we have rigorously proved the existence of the Born approximation.
@@
These experiments are presented in Section 4. First, we study the ability of the Born approximation $\gamma^{\mathrm{B}}_{1}$ to recover different bump perturbations of the conductivity $\sigma=1$ (which corresponds to $\kappa=0$). Second, we analyze how the resolution deteriorates with depth. Third, we consider conductivities that are perturbations of $\sigma_{\kappa,2}$ and analyze the difference between the Born approximations $\gamma^{\mathrm{B}}_{\sigma_{\kappa,2}}$ and $\gamma^{\mathrm{B}}_{\sigma_{0,2}}$. This experiment illustrates the significant advantage of using a good prior for $\sigma$. In fact, we show that the correct perturbation can be recovered from $\gamma^{\mathrm{B}}_{\sigma_{\kappa,2}}$ even in the presence of screening effects (low conductivity at the boundary), while $\gamma^{\mathrm{B}}_{\sigma_{0,2}}$ is completely useless for this task (see Fig. 5). Finally, we study how the method deals with discontinuous conductivities and noisy data in experiments 4 and 5, respectively.
1.2. Relation to other works in the literature

### v2

# Characterization of Dirichlet-to-Neumann maps via the Born approximation

@@ Abstract.
The problem of identifying the set of Dirichlet-to-Neumann (DtN) maps arising from conductivities on a smooth domain, among operators acting on functions on the boundary, is a challenging issue in the mathematical analysis of the Calderón inverse problem. This question is also relevant in specific applications since, as the inverse problem is ill-posed, numerically reconstructing a conductivity from the knowledge of its DtN map is particularly delicate. In this article, we address this issue by proving that any DtN map arising from a radial conductivity in the unit ball of $\mathbb{R}^{d}$ admits an exact representation as a linearized DtN map for a uniquely determined integrable function, that we call the Born approximation. This gives a strong necessary condition for an operator to be a DtN map arising from a radial conductivity. In particular, our results are a starting point towards developing a rigorous foundation for numerous linearization-based methods that are commonly used in the numerical solution of the Calderón inverse problem. We also characterize the Born approximation as a solution to a generalized moment problem that is formally well-defined even for non-radial conductivities. We investigate the uniqueness and structure of general non-radial solutions to this moment problem on the unit disk and provide an algorithm to numerically reconstruct the Born approximation in this setting. We provide numerical experiments to test the resolution and robustness of the Born approximation in different situations.
@@ 1.1. Setting and results
where $\gamma$ varies in some class of admissible conductivities. The mathematical analysis of this inverse problem has a long and rich history; results showing that $\gamma$ can be uniquely reconstructed from $\Lambda_{\gamma}$ (i.e., that $\Phi$ is injective) date back to the seminal works by Kohn and Vogelius [31], Sylvester and Uhlmann [42, 43], and Nachman [38]. Stability, another important aspect of the inverse problem, has also been widely studied. $\Phi^{-1}$ is known to be discontinuous for any natural topology on the set of DtN maps (the inverse problem is ill-posed) although numerous conditional stability results are known to hold. We refer the reader to the recent monograph [19] for a detailed description of these results as well as a thorough guide to the literature.
Much less is known about the problem of characterizing the range of $\Phi$, that is, the set of all DtN maps arising from conductivities (some partial results for very specific classes of two-dimensional conductivities can be found in [27, 40]); because of ill-posedness, this question is especially relevant to the reconstruction aspect of the Calderón problem: namely, that of finding effective procedures to construct $\gamma$ from $\Lambda_{\gamma}$. Reconstruction has been thoroughly investigated, partly due to its connection to a number of imaging techniques, such as Electrical Impedance Tomography (EIT).
The results in this article aim to contribute to the understanding of the range of $\Phi$, and as a byproduct, to provide a rigorous background for some widely used reconstruction methods in EIT based on linearization. A short review on these methods, and a description of their connection with the results in this article is presented in Section 1.2.
Linearization methods are based on the inversion of $\mathrm{d}\Phi_{\sigma}$, the Fréchet differential of $\Phi$ at some reference conductivity $\sigma$. This is known as the linearized Calderón problem. Since the work of Calderón [9], understanding this linearized version of the inverse problem has been a cornerstone in the strategy for solving the full nonlinear problem from both theoretical and numerical viewpoints. One tries to determine a function $\gamma^{\mathrm{B}}_{\sigma}$ that solves
@@
where $\gamma$ is the conductivity one desires to approximate with respect to a background conductivity $\sigma$ (or a sufficiently close prior); see [26] for instance. This is merely formal, since it is not clear that such a function exists. Motivated by the linear approximation in inverse scattering theory developed by Born in the 1920s, following [4, 5], we call $\gamma^{\mathrm{B}}_{\sigma}$ the Born approximation of $\gamma$ with respect to $\sigma$.
Though widely used in EIT, there are very few rigorous results on the existence of $\gamma^{\mathrm{B}}_{\sigma}$. Unlike in inverse scattering problems, the question of existence of the Born approximation in the Calderón problem is a subtle issue. If one were able to establish it, this would imply that every DtN map could be written as a linearized DtN operator, and therefore that the range of $\Phi$ is contained in that of $\mathrm{d}\Phi_{\sigma}$. To the best of our knowledge, the existence of $\gamma^{\mathrm{B}}_{\sigma}$ has been rigorously established only when $\sigma=1$, $\Omega=\mathbb{B}^{d}\subset\mathbb{R}^{d}$ is the unit ball, and $\gamma\in W^{2,p}(\mathbb{B}^{d})$ is radial; see [18, Theorem 1]111[18, Theorem 1] proves the existence of the Born approximation as the solution of a certain moment problem involving the spectrum of the DtN map. The connection with the Fréchet differential of $\Phi$ is presented in identity (2.27) in the proof of 1. (see also [17, 35] for rigorous results on the existence of the Born approximation for DtN maps associated with Schrödinger operators).
Furthermore, if $\gamma^{\mathrm{B}}_{\sigma}$ exists, then the strong injectivity properties of $\mathrm{d}\Phi_{\sigma}$ imply that one has a two-step factorization of the inverse problem: a linear step that asks for reconstructing the Born approximation $\gamma^{\mathrm{B}}_{\sigma}$ from the DtN map, and a non-linear one that consists of the inversion of the map $\Phi_{\sigma}^{\mathrm{B}}$ that sends the function $\gamma$ to the function $\gamma^{\mathrm{B}}_{\sigma}$. More precisely, using that $\Lambda_{\sigma}=\mathrm{d}\Phi_{\sigma}(\sigma)$ (see Proposition 2.1), (1.4) becomes
@@
This type of factorization can be formulated in a general context and has been applied under different names to solve a number of ill-posed inverse problems (see the survey [34]). In [18] it is shown that when $\gamma\in W^{2,p}(\mathbb{B}^{d})$ is a radial conductivity, $\gamma^{\mathrm{B}}_{1}$ exists and uniquely determines $\gamma$. Moreover, in this case, the map $(\Phi_{1}^{\mathrm{B}})^{-1}:\gamma^{\mathrm{B}}_{1}\longmapsto\gamma$ is Hölder continuous under suitable a priori assumptions on the conductivities (that unlike most conditional stability results, do not imply any compactness on the admissible set of conductivities). In other words, the aforementioned fact that $\Phi^{-1}:\Lambda_{\gamma}\longmapsto\gamma$ is discontinuous is due to the linear part of the factorization: the inversion of $\mathrm{d}\Phi_{\sigma}$. We say that the inverse problem admits a stable factorization if $\Phi$ can be factorized into a linear ill-posed map and a non-linear stable map. In addition to the results in [18] for the conductivity case, this property has also been proven in [17] for the radial Schrödinger case and an associated inverse spectral problem, and locally in [35] for the radial fixed energy Schrödinger case.
In this article, we present a systematic approach to establishing the existence of the Born approximation based on the well-known reduction to a problem involving Schrödinger operators; as a consequence of our analysis, we broaden the class of conductivities $\sigma$ and $\gamma$ for which (1.4) can be rigorously solved. More specifically, we show that $\gamma^{\mathrm{B}}_{\sigma}$ exists as an integrable function as soon as $\Omega=\mathbb{B}^{d}$ for $d\geq 2$, $\gamma$ is a $W^{2,\infty}$ radial conductivity, and $\sigma$ is
@@
We denote by $\mathcal{E}^{\prime}(\Omega)$ the space of distributions with compact support contained in $\Omega$. It is shown in Remark 2.2 that, for $\sigma\in\mathcal{C}^{\infty}(\overline{\Omega})$, one can extend $\mathrm{d}\Phi_{\sigma}$ from its natural domain $L^{\infty}(\Omega,\mathbb{R})$ to $L^{\infty}(\Omega)+\mathcal{E}^{\prime}(\Omega)$.
@@
Suppose that $\Omega:=\mathbb{B}^{d}$ is the unit ball in $\mathbb{R}^{d}$, that $\gamma\in W^{2,\infty}(\mathbb{B}^{d})$ is a radial conductivity, and that $\kappa\in(-\infty,\lambda_{\nu_{d},1}^{2})$. Then there exists a distribution $\gamma^{\mathrm{B}}_{\sigma_{\kappa,d}}\in W^{2,\infty}(\mathbb{B}^{d})+\mathcal{E}^{\prime}(\mathbb{B}^{d})$ such that:
(1.8) $\Lambda_{\gamma}=\mathrm{d}\Phi_{\sigma_{\kappa,d}}(\gamma^{\mathrm{B}}_{\sigma_{\kappa,d}}).$
In addition, the following statements hold:
@@
    $\gamma^{\mathrm{B}}_{\sigma_{\kappa,d}}$ is uniquely determined in $W^{2,\infty}(\mathbb{B}^{d})+\mathcal{E}^{\prime}(\mathbb{B}^{d})$ by (1.8).
  2. ii)
@@
1 generalizes to $\kappa\neq 0$ some of the results in [18]. The uniqueness statement actually holds in a slightly more general setting, see Theorems 2.7 and 2.8.
A second set of results aims to understand the validity of 1 for non-radial conductivities. We first observe that the Born approximation can be characterized as a solution to a certain complex moment problem that is described as follows. Let $(e_{m})_{m\in\mathbb{N}}$ be an orthonormal basis of $L^{2}(\partial\Omega)$ such that $e_{m}\in H^{1/2}(\partial\Omega)$ for all $m\in\mathbb{N}$. By direct computation of the Fréchet differential $\mathrm{d}\Phi_{\sigma}$ (see Proposition 2.1), one can prove that for all $F\in L^{\infty}(\Omega)$, and $\ell,m\in\mathbb{N}$
@@
$\mathfrak{m}_{\ell,m}^{\sigma}[F]:=\left\langle\overline{e_{\ell}},\mathrm{d}\Phi_{\sigma}(F)e_{m}\right\rangle_{H^{1/2}\times H^{-1/2}}=\int_{\Omega}F(x)\overline{\nabla u_{\ell}^{\sigma}}\cdot\nabla u_{m}^{\sigma}\,\mathrm{d}x,$
where $u_{m}^{\sigma}$ is the solution of (1.1) with $f=e_{m}$. By (1.5), it follows that $\gamma^{\mathrm{B}}_{\sigma}$ solves the generalized moment problem
(1.9) $\mathfrak{m}_{\ell,m}^{\sigma}[\gamma^{\mathrm{B}}_{\sigma}]=\int_{\Omega}\gamma^{\mathrm{B}}_{\sigma}\,\overline{\nabla u_{\ell}^{\sigma}}\cdot\nabla u_{m}^{\sigma}\,\mathrm{d}x=\left\langle\overline{e_{\ell}},\Lambda_{\gamma}e_{m}\right\rangle_{H^{1/2}\times H^{-1/2}},\qquad\ell,m\in\mathbb{N}.$
These remarks show that when $\Omega=\mathbb{D}$ is the unit disk in $\mathbb{R}^{2}$ and $\sigma=1$, the conclusion of 1 may fail for non-radial $\gamma$. In fact, letting $e_{m}(\theta):=(2\pi)^{-1/2}e^{im\theta}$ be the standard basis of $L^{2}(\mathbb{S}^{1})$ of spherical harmonics, we show (see Lemma 3.1) that for $F\in L^{1}(\mathbb{D})$,
$\mathfrak{m}_{\ell,m}^{1}[F]=0,\qquad\forall\ell,m\in\mathbb{Z},\;\ell m<0.$
As a consequence, $\mathrm{d}\Phi_{1}(F)$ maps $H^{1}(\mathbb{S}^{1})\cap\mathcal{H}_{+}$ into $\mathcal{H}_{+}$, where
$\mathcal{H}_{+}:=\{u\in L^{2}(\mathbb{S}^{1})\,:\,\left\langle e_{\ell},u\right\rangle_{L^{2}(\mathbb{S}^{1})}=0,\;\forall\ell<0\}.$
If the Born approximation $\gamma^{\mathrm{B}}_{1}$ were to exist, this would imply that $\Lambda_{\gamma}$ also maps $H^{1}(\mathbb{S}^{1})\cap\mathcal{H}_{+}$ into $\mathcal{H}_{+}$, which is not necessarily the case when $\gamma$ is non-radial. This can be checked in explicit examples: if $\varepsilon>0$ and $\gamma_{\varepsilon}(x,y):=1+\varepsilon(1-x^{2}-y^{2})(x^{2}-y^{2})$, with $(x,y)\in\mathbb{D}$, direct computation shows that:
$\left\langle\overline{e_{-3}},\Lambda_{\gamma_{\varepsilon}}e_{1}\right\rangle_{H^{1/2}\times H^{-1/2}}=-\frac{\varepsilon^{2}}{320}+\mathcal{O}(\varepsilon^{3}).$
Therefore, in this setting, it makes sense a priori to define the Born approximation $\gamma^{\mathrm{B}}_{1}$ as a function/distribution satisfying:
$\mathrm{d}\Phi_{1}(\gamma^{\mathrm{B}}_{1})=\Pi_{+}\Lambda_{\gamma}\Pi_{+}|_{H^{1}(\mathbb{S}^{1})},$
where $\Pi_{+}:L^{2}(\mathbb{S}^{1})\longrightarrow\mathcal{H}_{+}$ denotes the orthogonal projection. A characterization for $\gamma^{\mathrm{B}}_{1}$ based on this Ansatz, which turns out to be equivalent to solving the moment problem (1.9) with respect to the basis of spherical harmonics, is given in Section 3. More generally, the structure of $\gamma^{\mathrm{B}}_{\sigma}$ for $\sigma=\sigma_{\kappa,2}$ on $\mathbb{D}$ with $\kappa<\lambda_{0,1}^{2}$ is also investigated through the moment problem (1.9) (again with respect to the basis $(e_{m})_{m\in\mathbb{N}}$ of spherical harmonics in $\mathbb{S}^{1}$). In particular, (1.9) is shown to possess at most one solution in this case.
@@
We also numerically explore the structure and properties of the Born approximation as defined by (1.9) through several examples of non-radial $\gamma$, for $\sigma=\sigma_{\kappa,2}$ on the unit disk in $\mathbb{R}^{2}$. These experiments are presented in Section 4. This choice of $\sigma$ has the advantage of providing good a priori knowledge of the corresponding solutions $u_{m}^{\sigma}$ and encompasses the framework in which we have rigorously proved the existence of the Born approximation.
@@
First, we study the ability of the Born approximation $\gamma^{\mathrm{B}}_{1}$ to recover different bump perturbations of the conductivity $\sigma=1$ (which corresponds to $\kappa=0$). Second, we analyze how the resolution deteriorates with depth. Third, we consider conductivities that are perturbations of $\sigma_{\kappa,2}$ and analyze the difference between the Born approximations $\gamma^{\mathrm{B}}_{\sigma_{\kappa,2}}$ and $\gamma^{\mathrm{B}}_{\sigma_{0,2}}$. This experiment illustrates the significant advantage of using a good prior for $\sigma$. In fact, we show that the correct perturbation can be recovered from $\gamma^{\mathrm{B}}_{\sigma_{\kappa,2}}$ even in the presence of screening effects (low conductivity at the boundary), while $\gamma^{\mathrm{B}}_{\sigma_{0,2}}$ is completely useless for this task (see Fig. 5). Finally, we study how the method deals with discontinuous conductivities and noisy data in experiments 4 and 5, respectively.
1.2. The role of $\gamma^{\mathrm{B}}_{\sigma}$ in linearization methods for EIT